OpenStudy (anonymous):

log2 8 = 3 written in exponential form

rishavraj (rishavraj):

\[\log_a b = x \] \[b = a^x\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

2^3 = 8 2^8 = 3 3^2 = 8 3^8 = 2 but it's supposed to be one of these answers I think it's B but i'm not sure

rishavraj (rishavraj):

how could it be B....huh as i told u ......before..... have a look over tht stuff

rishavraj (rishavraj):

hav a look http://www.purplemath.com/modules/logs.htm

OpenStudy (anonymous):

C?

rishavraj (rishavraj):

how it could be tht ........????? http://prntscr.com/7syqe7